NvidiaIntely (nvidiaintely):

Find the height of a cylinder with radius 25 cm and surface area 8953.54 cm2. 32.00 cm 16.00 cm 31.00 cm 62.00 cm

NvidiaIntely (nvidiaintely):

@Google

OpenStudy (google):

You still got that site?

NvidiaIntely (nvidiaintely):

Which site?

NvidiaIntely (nvidiaintely):

yttechy.cf

OpenStudy (google):

yttechy

NvidiaIntely (nvidiaintely):

Yes I have it.

OpenStudy (google):

yeah why does it look like a 2004 site?

NvidiaIntely (nvidiaintely):

Because I like it that way for now. I will change it later :) Don't have time right now.

OpenStudy (google):

hmm ok

NvidiaIntely (nvidiaintely):

Can you help me with some questions?

OpenStudy (google):

I can help you with this one then i need to do something

NvidiaIntely (nvidiaintely):

Okay thanks!

NvidiaIntely (nvidiaintely):

@Google, are you going to start typing it?

OpenStudy (zeesbrat3):

Do you know the formula to find the surface area of a sphere?

OpenStudy (zeesbrat3):

\[SA = 2(\pi)rh + 2(\pi)r ^{2}\]

OpenStudy (zeesbrat3):

@NvidiaIntely

OpenStudy (zeesbrat3):

All you have to do is plug in your SA (Surface area) and your r (radius)

OpenStudy (zeesbrat3):

Therefore: \[8953.54 cm ^{2} = 2(\pi)(25 cm)(h) + 2(\pi)(25 cm)^{2}\]

OpenStudy (zeesbrat3):

\[8953.54cm ^{2} = 50(\pi)h + 1250(\pi) \]

OpenStudy (zeesbrat3):

\[5028.54 = 157h\]

OpenStudy (zeesbrat3):

\[h = 32.0289 cm\]

NvidiaIntely (nvidiaintely):

Thanks!

OpenStudy (zeesbrat3):

Any questions?

NvidiaIntely (nvidiaintely):

no